# Mean-Field Microcanonical Gradient Descent

### v1

@@ 4 Mean-Field Microcanonical Gradient Descent
The accuracy of the microcanonical approximation is based on the assumption of the concentration of the energy vector $\Phi(X)$ for the true distribution. If the microcanonical set is too large (i.e., too large of a value of $\epsilon$ or too small of a value of $T$), the microcanonical model would include many atypical outcomes that do not satisfy this concentration. On the other hand, the entropy of the microcanonical distribution is the log of the volume of the microcanonical set $\Omega_{\varepsilon}$, and is therefore increasing in the radius $\varepsilon$. Thus, having too small of a radius may ensure no atypical samples are generated, but at the price of a loss of entropy, and will therefore result in a larger KL divergence.
Given that the energy function is appropriate for the distribution, e.g. it is taken to be the sufficient statistics of the distribution as in the AR(1) case, the likelihood increases as the descent progresses and the energy vector approaches the target energy. Conversely, if too many iterates are performed, the energy vectors of the samples will be too close, effectively choosing a radius that is too small.
@@
Figure 2: Illustration of $\Phi$-pushforward measures of the true distribution in blue centered at the origin, and the approximation in orange. In the regular MGDM, each particle individually seeks to minimize its distance to the origin in energy space, potentially causing a collapse; in the mean-field version, the particles move approximately in parallel in energy space.
@@
This observation leads to our proposition of the mean-field microcanonical gradient descent model (MF–MGDM), where the mass of the initial distribution is pushed towards the target in energy space while attempting to reduce the collapse of the radius of the ball (or similarly the energy variance) and thereby reducing the entropy loss. The principle is illustrated in Figure 2; whereas the regular MGDM (Figure 2(a)) updates each sample individually with the objective of minimizing its energy distance (2) to the target, MF-MGDM (Figure 2(b)) updates several samples simultaneously so that they move towards the target energy in aggregate.
@@
The mean-field concept originates from statistical physics as a tool for studying macroscopic phenomena in large particle systems by averaging over microscopic interactions. In the context of game theory, mean-field games are multiagent problems where each agent has a negligible impact on the others, so that the dynamics of an agent depends on the law of the system. For an $N$-player system, the law is the empirical measure, for which a subclass of systems are those where the dynamics depend on the empirical mean. The mean-field limit is then when $N\to\infty$; see e.g. (Carmona & Delarue, 2018).
A heuristic derivation of the MF–MGDM is the following: Define $\bm{x}=\{x^{(n)}\}_{n=1}^{N}\in\mathcal{X}^{N}$ as a collection of $N$ particles, whereby particle we mean a sample path, and denote the batch mean energy as
$\bar{\Phi}(\bm{x})=\frac{1}{N}\sum_{i=1}^{N}\Phi(x^{(i)}).$
We would like to update each particle so that the particles in aggregate moves towards the target energy, i.e. the origin, but requiring that they move in the same direction in energy space. Define this direction as
$h(\bm{x})=-\gamma\bar{\Phi}(\bm{x}),$
emulating a force on the batch mean towards the target as illustrated in Figure 2(b). We want to update $\bm{x}$ to $\bm{x}^{\prime}$ so that $\Phi(\bm{x}^{\prime})\approx\Phi(\bm{x})+h(\bm{x})$, which can be seen as an inner optimization problem to minimize the objective $\frac{1}{2}\|\Phi(\bm{x}^{\prime})-(\Phi(\bm{x})+h(\bm{x}))\|$ where we can let $\bm{\xi}_{0}=\bm{x}$ and iteratively perform the update
$\bm{\xi}_{\tau+1}$ $=\bm{\xi}_{\tau}-\frac{\lambda}{2}\nabla_{\bm{\xi}}\|\Phi(\bm{\xi}_{\tau})-(\Phi(\bm{x})+h(\bm{x}))\|^{2}$
$=\bm{\xi}_{\tau}-\lambda J_{\Phi}^{\top}(\bm{\xi}_{\tau})\left[\Phi(\bm{\xi}_{\tau})-(\Phi(\bm{x})+h(\bm{x}))\right]$
for some number of iterations, where $\lambda$ is the inner step size. Taking only one step in this procedure results in the update rule
$\bm{x}^{\prime}=\bm{\xi}_{1}$ $=\bm{x}-\lambda J_{\Phi}^{\top}(\bm{x})\left[\Phi(\bm{x})-(\Phi(\bm{x})+h(\bm{x}))\right]$
$=\bm{x}-\lambda\gamma J_{\Phi}^{\top}(\bm{x})\bar{\Phi}(\bm{x}).$
Redefining $\gamma$ as the composite of step lengths in both internal and external optimization, we define the mean-field gradient step as
$\tilde{g}(\bm{x})=\bm{x}-\gamma J_{\Phi}^{\top}(\bm{x})\bar{\Phi}(\bm{x}).$ (6)
It is possible to take more steps in the inner iteration, although this would increase the number of times the Jacobian of $\Phi$ has to be computed. From a sampling perspective, it can still be feasible depending on the energy function, since the computation is quite fast, but evaluating the KL divergence can become too costly. Empirically, it seems that one inner iteration is sufficient for the purpose of avoiding collapse, so for the sake of being able to compute KL efficiently, we restrict the model to only one inner iteration.
@@
The MF–MGDM faces two challenges that the regular model does not. The first is that the sampling procedure requires simultaneous generation of multiple samples in order to compute $\bar{\Phi}$. This is solved efficiently by vectorizing the computation of the energy Jacobian in (6). The second challenge is that of computing the entropy, specifically computing the log-determinant of the Jacobian of a gradient step $\tilde{g}$. The issue is that the samples are now coupled, resulting in the Jacobian being one large $Nd\times Nd$ matrix, where the off-diagonal $d\times d$ blocks are non-zero. Naively computing the determinant scales as $\operatorname{\mathcal{O}}(N^{3}d^{3})$ (even keeping it in memory is infeasible), but it is possible to rewrite it on a form that allows $\operatorname{\mathcal{O}}(Nd^{3})$ computation by writing the Jacobian as a sum of a block diagonal and a low-rank matrix, and then using the matrix determinant lemma (see Appendix A).
(a) Energies after 157 steps
(b) Reverse KL Divergence
Figure 3: MF-MGDM for AR(1) example.
@@
To demonstrate the effect of the mean-field gradient step, we return to the AR(1) example. Figure 3(a) shows the pushforward by $\Phi$ of the initial measure, the final measure after 157 steps, and the true measure. We see now that the final distribution in energy space more closely aligns with that of the true measure, preventing the reduction of entropy observed in the MGDM (see Figure 1(a)). Tracking the reverse KL divergence for each gradient step, Figure 3(b), we see a monotone decrease, avoiding the need for early stopping. If we break up the KL divergence into negative entropy and log-likelihood, we no longer observe an unbounded decrease in entropy. Instead, it stabilizes around a value close to the negative log-likelihood, resulting in a small KL divergence.
In Figure 4, the impact of the choice of $N$ is illustrated. For larger $N$, the estimate $\bar{\Phi}$ of $\mathbb{E}_{q}[\Phi(X)]$ improves, the model also improves in terms of reverse KL divergence. It is also apparent that for $N$ large enough, there are only marginal improvements.
@@
Figure 4: Reverse KL divergence through gradient descent with respect to the true model AR(1) for MF–MGDM with different mean-field batch sizes. The KL estimates are computed from 128 samples.
It is worth pointing out that, although the MF–MGDM is an attempt at maintaining the variance in energy from the initial distribution, this is not guaranteed. For example, with $\Phi(x)\coloneqq\|x\|_{2}^{2}$, we have the update
$x_{t+1}=x_{t}-2\gamma x_{t}\bar{\Phi}(\bm{x}_{t})=(1-2\gamma\bar{\Phi}(\bm{x}_{t}))x_{t}$
for an arbitrary particle $x$, so in the mean-field limit,
$\Phi(x_{t+1})=(1-2\gamma\mathbb{E}_{\mu_{t}}(\Phi(X)))^{2}\Phi(x_{t})$
and, provided $\mathbb{E}_{\mu_{0}}(\Phi(X))\neq 0$, we obtain
$\operatorname{Var}_{\mu_{t+1}}(\Phi(X))$ $=(1-2\gamma\mathbb{E}_{\mu_{t}}(\Phi(X)))^{4}\operatorname{Var}_{\mu_{t}}(\Phi(X))$
$<\operatorname{Var}_{\mu_{t}}(\Phi(X)).$
@@ 5 Numerical Experiments
Table 1: Reverse KL divergence for different distributions and approximation models
@@
AR$(0.1)$ 2.76 0.09 4.24 1.99 5.47 4.01 5.44 2.32
AR$(0.2,-0.1)$ 9.44 3.81 17.98 10.55 25.91 14.84 27.33 9.78
AR$(-0.1,0.2,0.1)$ 30.04 26.39 20.98 15.18 29.55 18.01 28.46 10.03
CIR$(2^{-1},1,1)$ 219.40 214.65 170.63 169.56 121.17 59.21 105.05 30.78
CIR$(2^{-1/2},2^{1/2},1)$ 104.49 87.96 182.24 179.52 223.63 204.79 203.44 201.42
5.1 Synthetic data
To compare the different approximation models on synthetic data, we use time series models that have density functions in closed form, allowing for evaluation of the reverse KL divergence. We generate 10 000 samples of length $1024$ and take the average energy over these samples as target energy, to simulate the idealized setting where the true energy vector is known, avoiding bias. The KL divergence is estimated by generating 128 samples from the respective models, recording the divergence after each gradient step.
The energy functions used are
@@ 5.1 Synthetic data
    mean (over time) of second-order scattering transform (with complex modulus as nonlinearity), using filters from the Kymatio package (Andreux et al., 2020);
@@
    covariance (over time) of second-order scattering transform, augmented with filters shifted by 00 and $\pi/3$ in the first-order coefficients, and using ReLU of the real part as nonlinearity. Finally, we perform a dimensionality reduction by using principal component analysis (PCA) on transforms applied to Gaussian white noise;
@@
    scattering spectra from (Morel et al., 2023), taking the modulus of those coefficients which are complex, and thereby ignoring the phase.
@@
An $\text{AR}(p)$ model with parameters $\varphi_{1},\dots,\varphi_{p}$ and $\sigma$ is a generalization of the AR(1) process in (4) and is defined by the recursion
$x_{i}=\sum_{j=1}^{p}\varphi_{j}x_{i-j}+\sigma\varepsilon_{i},$
with white noise $(\varepsilon_{i})_{i}$, and is stationary if the roots of the characteristic polynomial $\pi(z)=1-\sum_{j}\varphi_{j}z^{j}$ are outside the unit circle. In Table 1, the models are denoted AR$(\gamma_{1},\gamma_{2},\dots)$, and $\sigma$ is chosen as to get unit marginal variance.
@@
The CIR model (Cox et al., 1985) is a diffusion process that is commonly used for modelling short-term interest rates. It is related to the Ornstein–Uhlenbeck process, which can be seen as a continuous version of AR(1), but differs in the way that the diffusion term is scaled by the square root of the rate $r\in\mathbb{R}^{+}$:
@@
where $W$ is a Brownian motion. The process admits a stationary distribution, and the distribution at time $t$ given the value at an earlier time $s<t$ can be written in closed form, allowing for explicit evaluation of the likelihood of a discretization in an autoregressive fashion $\prod_{i}p(r_{i\Delta t}\mid r_{(i-1)\Delta t})$. The distribution of $r_{0}$ can be taken to be the marginal distribution, i.e., a gamma distribution. In the experiments, we let $\Delta t=1$, and the models are identified as CIR$(\kappa,\theta,\sigma)$.
The CIR process is non-negative, so projected gradient descent, described in Appendix B, has to be used when approximating this distribution in the context of MGDM. In addition, the support of the distribution is $\mathbb{R}^{+}$, so given that it exists, the maximum entropy distribution conditioned on the first two moments is the exponential if the mean and standard deviation are equal, otherwise a truncated Gaussian (Dowson & Wragg, 1973).
@@ Results
For each model and energy function, the reverse KL divergence was computed at each step through the descent. The value at the step where minimum divergence was achieved is displayed in Table 1. For every true distribution, we present results also for energy functions that are not necessarily a good choice, given the true model. This is to point out that the MF–MGDM consistently performs at least as well as the MGDM.
The behavior of the divergence through the descent, as well as its constituents entropy and expected log-likelihood, are displayed in Figure 5, using the energy function that best approximates each distribution in accordance with Table 1. Here we see again that less entropy is lost in the mean-field model, and that there is a marginal difference in general between the expected likelihoods of the two models.
(a) AR$(0.1)$ with ACF as energy.
(b) AR$(0.2,-0.1)$ with ACF as energy.
(c) AR$(-0.1,0.2,0.1)$ with scattering spectra as energy.
(d) CIR$(2^{-1},1,1)$ with scattering spectra as energy.
(e) CIR$(2^{-1/2},2^{1/2},1)$ with ACF as energy.
Figure 5: Reverse KL divergence (left), negative entropy (right, solid) and log-likelihood (right, dashed) through the descent. The energy function used for each distribution is the corresponding optimal energy function according to Table 1.
Figure 6: S&P 500 autocorrelation, autocorrelation of the square of the signal, and marginal histogram. The top two rows are violin plots illustrating the empirical marginal density of the statistics for the models, with whiskers indicating min and max. The statistics of the true distribution are in black in all rows. In the bottom row, the dashed gray line is a Gaussian with moments matched to the true data.
Another important difference here is that while MGDM needs to be stopped early to prevent the entropy from collapsing, this is not the case for MF–MGDM. Indeed, we see that the entropy stabilizes after a certain number of steps similarly to the log-likelihood. This is important because in a real-world setting, the true distribution is not known and the reverse KL divergence is not computable, so we cannot reasonably estimate the number of gradient steps to perform in order to balance the entropy loss with the increase of expected log-likelihood. For MF–MGDM, we can run the sampling until convergence without having to worry about this type of overfitting.

### v2

@@ 4 Mean-field microcanonical gradient descent
In the MGDM, the expected log-likelihood increases as the descent progresses and the energy approaches the target (assuming an appropriate energy function for the given distribution we model, e.g. the sufficient statistics as in the AR(1) case). Conversely, if too many iterates are performed, the energy vectors of the samples will be too close. Note that this happens even if the $\varepsilon$ parameter of the microcanonical ensemble is chosen to be large, since the MGDM method will be concentrated over a small subset of $\Omega_{\varepsilon}$. This observation leads to our proposition of the mean-field microcanonical gradient descent model (MF–MGDM).
4.1 The model
@@
Figure 3: Illustration of $\Phi$-pushforward measures of the true distribution in blue centered close to the target energy $\alpha$, and the approximation in orange. In the regular MGDM, each particle individually seeks to minimize its distance to the origin in energy space, potentially causing a collapse; in the mean-field version, the particles move approximately in parallel.
@@
In the MF–MGDM, the mass of the initial distribution is pushed towards the target in energy space while attempting to reduce the collapse of the radius of the ball (or similarly the energy variance) and thereby reducing the entropy loss. The principle is illustrated in Figure 3. Whereas the regular MGDM (Figure 3(a)) updates each sample individually with the objective of minimizing its energy distance (3) to the target, MF–MGDM (Figure 3(b)) updates several samples simultaneously so that they move towards the target energy in aggregate.
Formally, define $\bm{x}=\{x^{(n)}\}_{n=1}^{N}\in\mathbb{R}^{Nd}$ as a collection of $N$ particles, where a particle is a sample path in $\mathbb{R}^{d}$. Recalling the mean energy $\overline{\Phi}$ in (5), the new optimization objective is
$\overline{L}(\bm{x})\coloneqq\frac{N}{2}\|\overline{\Phi}(\bm{x})-\alpha\|^{2}.$ (9)
Denoting by $\mathcal{J}_{\Phi}(\bm{x})$ the concatenation of the Jacobians $J_{\Phi}(x^{(n)})$ of $\Phi$ with respect to each particle $x^{(n)}$,
$\mathcal{J}_{\Phi}(\bm{x})\coloneqq\begin{bmatrix}J_{\Phi}(x^{(1)})&\cdots&J_{\Phi}(x^{(N)})\end{bmatrix}\in\mathbb{R}^{K\times Nd},$ (10)
we define the mean-field gradient step as a gradient step for the objective (9), namely
$\overline{g}(\bm{x})\coloneqq\bm{x}-\gamma\mathcal{J}_{\Phi}^{\top}(\bm{x})\left(\overline{\Phi}(\bm{x})-\alpha\right).$ (11)
@@
The mean-field concept originates from statistical physics as a tool for studying macroscopic phenomena in large particle systems by averaging over microscopic interactions. In the context of game theory, for instance, mean-field games are multiagent problems where each agent has a negligible impact on the others, so that the dynamics of an agent depends on the law of the system. For an $N$-player system, the law is the empirical measure, for which a subclass of systems are those where the dynamics depend on the empirical mean. The mean-field limit is then when $N\to\infty$; see e.g. Carmona and Delarue [30]. We can think of (11) as corresponding to a discretization of a system of differential equations with mean-field interactions.
The MF–MGDM faces two challenges that the regular model does not. The first is that the sampling procedure requires simultaneous generation of multiple samples in order to compute $\overline{\Phi}$. This is solved efficiently by vectorizing the computation of $\mathcal{J}_{\Phi}(\bm{x})$ in (10). Furthermore, most applications call for generation of multiple samples, so the additional cost would be incurred at any rate by multiple invocations of MGDM.
@@
The second challenge is that of computing the entropy, specifically computing the log-determinant of the Jacobian of a gradient step $\overline{g}$. The issue is that the samples are now coupled, resulting in the Jacobian being one large $Nd\times Nd$ matrix. Naively computing the determinant scales as $\operatorname{\mathcal{O}}(N^{3}d^{3})$ (even keeping the Jacobian in memory is infeasible), but it is possible to rewrite it on a form that allows $\operatorname{\mathcal{O}}(Nd^{3})$ computation by writing the Jacobian as a sum of a block diagonal and a low-rank matrix, and then using the matrix determinant lemma (see Appendix A).
4.2 An illustrative example – revisited
@@
To demonstrate the effect of the mean-field gradient step, we return to the AR(1) example. Figure 1(c) shows the pushforward by $\Phi$ of the MF–MGDM approximation after 157 steps when minimum KL is achieved. We see now that the final distribution in energy space more closely aligns with that of the true measure, preventing the reduction of entropy observed in the MGDM (see Figure 1(a)). Tracking the reverse KL divergence for each gradient step, Figure 2(b), we see an almost monotone decrease, avoiding the need for early stopping. If we break up the KL divergence into negative entropy and log-likelihood, we no longer observe an unbounded decrease in entropy. Instead, it stabilizes around a value close to the negative log-likelihood, resulting in a small KL divergence.
@@
Figure 4: Reverse KL divergence through gradient descent with respect to the true model AR(1) for MF–MGDM with different mean-field batch sizes $N$, and with Monte Carlo sample size 128.
4.3 Theoretical entropy bound
Since the entropy of $d$ i.i.d. random variables scales linearly in $d$ it is natural to define the entropy rate $d^{-1}H(p_{1}\times\cdots\times p_{d})$ for the joint distribution of sequences of random variables [4]. In MF–MGDM, the joint distribution is over $N$ time series of length $d$, hence we have to normalize with $Nd$.
Theorem 4.1.
Assume $\Phi\in\bm{\mathrm{C}}^{2}$, with $\beta$ and $\eta$ denoting the Lipschitz constants of $\Phi$ and $\nabla\Phi$, respectively. Denote $\overline{q}_{T}^{N}$ as the distribution of the MF–MGDM model with $N$ particles after $T$ iterations. Then the entropy rate $(Nd)^{-1}H(\overline{q}_{T}^{N})$ admits, up to $\operatorname{\mathcal{O}}(\gamma^{2})$ terms, the lower bound
$(Nd)^{-1}H(\overline{q}_{T}^{N})\geq(Nd)^{-1}H(\overline{q}_{0}^{N})-2\gamma\left(\eta\sqrt{K}\sum_{t=0}^{T-1}\mathbb{E}_{\overline{q}_{t}^{N}}\|\overline{\Phi}(\bm{X})-\alpha\|+\frac{K}{Nd}\beta^{2}T\right).$
The entropy bound for the regular MGDM is recovered when $N=1$ (since $\overline{\Phi}$ and $\Phi$ are then equal). Herein lies an explanation for the improvement in KL of the MF–MGDM. In both models, $\overline{\Phi}$ or $\Phi$ goes to $\alpha$, whereby the cost in entropy for each gradient step is after a point mainly driven by the $\beta^{2}T$-term, which can be made arbitrarily small in MF–MGDM by increasing $N$. This is also reflected empirically in Figure 4 where a monotonic improvement of KL divergence is observed as $N$ grows. Note, however, that this is a lower bound, so it does not guarantee that MF–MGDM always preserves entropy better than MGDM (although this is observed numerically), but it does provide a better guarantee. The proof of Theorem 4.1 is given in Appendix B.
@@ 5 Numerical experiments
5.1 Synthetic data
To compare the different approximation models on synthetic data, we use time series models that have density functions in closed form, allowing for evaluation of the reverse KL divergence. We generate 10 000 samples of length 1 024 and take the average energy over these samples as target energy, to simulate the idealized setting where the true energy vector is known, avoiding bias. The KL divergence is estimated by generating 128 samples from the respective models and recording the divergence after each gradient step.
Table 1: Minimum reverse KL divergence over $T$ for different distributions and approximation models, where Reg. denotes the regular MGDM whereas MF is the mean-field version; $N=128$.
@@ 5.1 Synthetic data
AR$(0.1)$ 2.76 0.09 4.24 1.99 5.47 4.04 5.44 2.32
AR$(0.2,-0.1)$ 9.44 3.81 17.98 10.55 25.91 14.84 27.33 9.60
AR$(-0.1,0.2,0.1)$ 30.04 26.39 20.98 15.18 29.55 18.01 28.46 10.13
CIR$(1/2,1,1)$ 219.40 214.65 170.99 168.88 121.17 59.21 105.05 30.78
CIR$(1/\sqrt{2},\sqrt{2},1)$ 104.49 87.96 182.32 179.34 223.63 204.79 203.46 201.44
We used the following energy functions:
@@
    First moments of the second-order scattering transform (with complex modulus as nonlinearity), using filters from the Kymatio package [31, v0.3] (BSD-3);
@@
    Second moments of the second-order scattering transform, augmented with filters shifted by 00 and $\pi/3$ in the first-order coefficients, and using ReLU of the real part as nonlinearity. Finally, we perform a dimensionality reduction by using principal component analysis (PCA) on transforms applied to Gaussian white noise;
@@
    Scattering spectra from Morel et al. [10] (MIT License), taking the modulus of those coefficients which are complex, and thereby ignoring the phase.
@@
An $\text{AR}(p)$ model with parameters $\varphi_{1},\dots,\varphi_{p}$ and $\sigma$ is a generalization of the AR(1) process in (6) and is defined by the recursion $x_{i}=\sum_{j=1}^{p}\varphi_{j}x_{i-j}+\sigma\varepsilon_{i},$ with white noise $(\varepsilon_{i})_{i}$, and is stationary if the roots of the characteristic polynomial $\pi(z)=1-\sum_{j}\varphi_{j}z^{j}$ are outside the unit circle. In Table 1, the models are denoted AR$(\varphi_{1},\dots,\varphi_{p})$, and $\sigma$ is chosen as to obtain unit marginal variance.
@@
The CIR model [32] is a diffusion process that is commonly used for modelling short-term interest rates. It is related to the Ornstein–Uhlenbeck process, which can be seen as a continuous version of AR(1), but differs in the way that the diffusion term is scaled by the square root of the rate $r\in\mathbb{R}^{+}$ to give
@@
where $W$ is a Brownian motion. The process admits a stationary distribution, and the distribution at time $t$ given the value at an earlier time $s<t$ is a scaled noncentral $\chi^{2}$ distribution which can be written in closed form, allowing for explicit evaluation of the likelihood of a discretization in an autoregressive fashion. The distribution of $r_{0}$ can be taken to be the marginal distribution, i.e., a gamma distribution. In the experiments, we use the discretization $\Delta t=1$, and the models are identified as CIR$(\kappa,\theta,\sigma)$. The CIR process is non-negative, so projected gradient descent, described in Appendix C, has to be used when approximating this distribution in the context of MGDM.
@@ Results
For each model and energy function, the reverse KL divergence was computed at each step through the descent. The minimum divergence achieved is displayed in Table 1. For every true distribution, we present results also for energy functions that are not necessarily a good choice, given the true model. We see here that MF–MGDM consistently outperforms MGDM.
The KL divergence through the descent as a function of iteration number is shown in Figure 5, as well as its constituents entropy and expected log-likelihood. Here we have only plotted results for the energy function that best approximates each distribution in accordance with Table 1. Here we see again that the mean-field model retains more entropy, and the difference is marginal between the expected likelihoods of the two models.
(a) AR$(0.2,-0.1)$
(b) AR$(-0.1,0.2,0.1)$
(c) CIR$(1/2,1,1)$
(d) CIR$(1/\sqrt{2},\sqrt{2},1)$
Figure 5: Reverse KL divergence (top), negative entropy (bottom, solid) and log-likelihood (bottom, dashed) through the descent. Blue is regular MGDM and orange is MF–MGDM. The energy function used for each distribution is the corresponding optimal energy function according to Table 1, i.e., (a) and (d) use ACF while (b) and (c) use scattering spectra. $N=128$.
Another important difference here is that while MGDM needs to be stopped early to prevent the entropy from collapsing, this is not the case for MF–MGDM. Indeed, we see that the entropy stabilizes after a certain number of steps similarly to the log-likelihood. This is important because in a real-world setting, the true distribution is not known and the reverse KL divergence is not computable, so we cannot reasonably estimate the number of gradient steps to perform in order to balance the entropy loss with the increase of expected log-likelihood. For MF–MGDM, we can run the sampling until convergence while being less sensitive to this type of overfitting.